# From Logits to Hierarchies: Hierarchical Clustering made Simple

### v1

@@ Appendix C Additional visualizations
In Figures 7 we provide additional visualizations for the hierarchies obtained with our proposed method in our hierarchical clustering experiments, complementing the quantitative and qualitative evidence shown in Section 4.1. Leafs are matched to the original labels by checking the most frequent label among datapoints contained in the leaf. In addition to the matched label, we report purity of each leaf, in percent.
@@
In Figure 8, we provide additional results for ImageNet with different colorings for the inferred hierarchy, supplementing our results from Section 4.2. These visualizations show where the subtree of birds (used in Figure 3(b)) is located within the complete tree and in relation to other superclasses, such as mammals, reptiles, dogs, and clothing.
@@
Figure 8: Visualization of the hierarchical clustering produced by L2H-TURTLE for the ImageNet-1K dataset. We show the complete tree of 1K classes colored by the corresponding WordNet hypernyms “bird”, “dog”, and “clothing” (Figure 8(a)) and by “bird”, “mammal”, and “reptile” (Figure 8(b)).

### v2

@@ Appendix C Additional Results and Visualizations
In Figures 6, 7, 8, we provide additional visualizations for the hierarchies obtained with our proposed method in our hierarchical clustering experiments, complementing the quantitative and qualitative evidence shown in Section 4.1. Note that, to produce these visualizations, as well as the other visualizations of hierarchies in this paper we used the iTOL tool (itolpaper). Leaves are matched to the original labels by checking the most frequent label among data points contained in the leaf. In addition to the matched label, we report the purity of each leaf in percentage.
@@
In Figure 9, we provide additional results for the ImageNet case study (Section 4.2) with different colorings for the inferred hierarchy, supplementing our results from Section 4.2. These visualizations show where the subtree of birds (used in Figure 3(b)) is located within the complete tree and in relation to other superclasses, such as mammals, reptiles, dogs, and clothing.
@@
Figure 9: Visualization of the hierarchy produced by our method in the experiment on ImageNet from Section 4.2. We show the complete tree of 1k classes colored by the corresponding WordNet hypernyms “bird”, “dog”, and “clothing” (Figure 9(a)) and by “bird”, “mammal”, and “reptile” (Figure 9(b)).